OpenStudy (liizzyliizz):

find the constants a and b such that the function is continuous on the entire real number line? What does it mean by "constants" in this case?

OpenStudy (liizzyliizz):

ill draw out the function in this case

OpenStudy (anonymous):

function is?

OpenStudy (anonymous):

is it given peicewise?

OpenStudy (anonymous):

if so plug in the values at the endpoints and solve for a and b

OpenStudy (liizzyliizz):

yess lol

OpenStudy (liizzyliizz):

oh ok.

OpenStudy (anonymous):

so say it changes at 2 replace x by 2, get an expression in a and b then if it changes again at 3 replace x by 3 get another expression in a and b solve the system

OpenStudy (anonymous):

go ahead and post, we can work it out if what i wrote is not clear

OpenStudy (liizzyliizz):

im trying to write the function but its coming out weird on here today. its basically x^3 if x is less than or equal to 2, and ax^2 if x > 2

OpenStudy (liizzyliizz):

in this case you would just plug in 2?

OpenStudy (anonymous):

ok so right away we want \[ 2^3=a\times 2^2\]

OpenStudy (anonymous):

just plug in 2 for x. solve for with your eyeballs, get \[a=2\]

OpenStudy (anonymous):

that way the functions will match up at 2. the limit will be 8 from the left and from the right.

OpenStudy (anonymous):

then if it changes again at some other place, repeat the process

OpenStudy (liizzyliizz):

so if lets say there are 3 parts to the piecewise, all 3 would have to equal to eachother?

OpenStudy (anonymous):

well there can't be three parts at one place. there can only be two right? but it can be defined differently at different places, so just plug in the value of x where it changes places into both formulas, set them equal and solve

OpenStudy (anonymous):

i saw one like this were you had to actually solve a system you don't in this case because you know a = 2 right away

OpenStudy (liizzyliizz):

f(x)= 2 x\[\le\]-1 ax+b -1<x<3 -2 x\[\ge\] 3

OpenStudy (liizzyliizz):

ok this doesnt like me very well, but i tried to write the function -.-

OpenStudy (anonymous):

now you get a system i think. replace x by -1 get \[-2 = -a+b\]

OpenStudy (anonymous):

i replaced x by -1 in top two. then replace x by 3 get \[3a+b = -\]

OpenStudy (anonymous):

dang \[3a+b=-6\]

OpenStudy (anonymous):

so solve the system \[-a+b=-2\] \[3a+b=-6\] and you will have your a and your b

OpenStudy (anonymous):

i gotta run but i hope it is clear what i did. make the replacement for x in both, got two equations for a and b. then solve

OpenStudy (liizzyliizz):

okey dokes thank you. I'll try solving this the best i can